OpenStudy (anonymous):

Simplify the expression four root of order three of x times x to the power of start fraction three over two end fraction end power times y to the power of six end power times y to the power of start fraction one over two end fraction end power and rewrite it in radical form.

OpenStudy (misty1212):

\[\huge \sqrt[4]{x^3}\times x^{\frac{3}{2}}\]?

OpenStudy (misty1212):

and \[\huge y^6\times y^{\frac{1}{2}}\]?

OpenStudy (misty1212):

\[y^6\times y^{\frac{1}{2}}=y^{6+\frac{1}{2}}=y^{\frac{13}{2}}\]

OpenStudy (anonymous):

\[4\sqrt[3]{x}*x^\frac{ 3 }{ 2 }*y^6*y ^\frac{ 1 }{ 2 }\]

OpenStudy (misty1212):

\[4x^{\frac{1}{3}+\frac{3}{2}}=4x^{\frac{11}{6}}\]

OpenStudy (misty1212):

so \[4\sqrt[6]{x^{11}}\times \sqrt{y^{13}}\]

OpenStudy (anonymous):

i got \[4\sqrt[11]{x^6} \sqrt[13]{y^2}\]

OpenStudy (anonymous):

is that correct?